### v1

# DiVE-k: differential visual reasoning for fine-grained image recognition

Abstract
Large Vision Language Models (LVLMs) possess extensive text knowledge but struggles to utilize this knowledge for fine-grained image recognition, often failing to differentiate between visually similar categories. Existing fine-tuning methods using Reinforcement Learning (RL) with exact string-match reward are often brittle, encourage memorization of training categories, and fail to elicit differential reasoning needed for generalization to unseen classes. To address this, we propose DiVE-k, Differential Visual rEasoning using top-k generations, framework that leverages model’s own top-k predictions as a training signal. For each training image, DiVE-k creates a multiple-choice question from the model’s top-k outputs and uses RL to train the model to select the correct answer. This approach requires the model to perform fine-grained differential reasoning among plausible options and provides a simple, verifiable reward signal that mitigates memorization and improves generalization. Experiments on five standard fine-grained datasets show that our method significantly outperforms existing approaches. In the standard base-to-novel generalization setting, DiVE-k surpasses the QWEN2.5-VL-7B and ViRFT by $10.04\%$ and $6.16\%$ on the Harmonic Mean metric, respectively. Further experiments show similar gains in mixed-domain and few-shot scenarios.  [^1] Our code is available at https://github.com/raja-kumar/DiVE-k
1 Introduction
We explore the task of zero-shot fine-grained image recognition, as a visual reasoning task building on available Large Vision Language Models (LVLMs). Such capabilities are crucial for the generalization of vision systems. In early zero-shot image recognition works, such as in CLIP (radford2021learning), the visual embedding from an image is matched against the text embedding of class names to determine the most likely label. LVLMs, such as QWEN2-VL (wang2024qwen2), contain a Large Language Model (LLM) in themselves and are able to use their vast language knowledge with unified multimodal pre-training to achieve impressive capabilities in zero-shot recognition. however, the accuracy for fine-grained recognition is limited. We aim to improve accuracy by fine-tuning on a subset of categories of a new dataset (called the “base” set) and test on “novel” categories for which no training examples are seen. This setting is common (zhou2022learning) and relevant for adapting models to new domains with limited training data.
Our approach is based on two key observations. First is that, the base model exhibits high variance across its Pass@K performance: the correct label often appears among the $K$ sampled response, yet fails to get it correct as Pass@1, see Figure 1 (a). This indicates possible over-reliance on coarse, salient attributes shared by related categories and may benefit from a fine-grained, differential reasoning to separate semantically similar categories. The second observation is that the LVLMs actually contain detailed knowledge about the parts and attributes of the base and novel categories which could be used for detailed, differential analysis.
Inspired by the success of Chain of Thought (CoT) reasoning (wei2022chain) using Reinforcement Learning (RL) (guo2025deepseek) for pure language tasks such as mathematics and coding (shao2024deepseekmath; jiang2023mistral7b), ViRFT (liu2025visual) extended this idea to vision tasks such as image classification. The key idea in ViRFT is to construct a verifiable reward (lambert2024tulu) for image classification which allows visual reasoning RL training via GRPO (shao2024deepseekmath). However, their verifiable reward obtained through exact string match between the category name and the model’s final answer is brittle: (i) requiring ad hoc string post-processing and model responses may use scientific or common names that cannot be validated by string matching; (ii) encourages memorization of the training category names (Section 4.2.1); and (iii) fails to incentivize attribute-level, discriminative reasoning see Figure 1b. This leads to a weak base to novel generalization and a tendency to ignore useful text knowledge when visual evidences alone are ambiguous.
To overcome these deficiencies, we propose DiVE-k framework (Differential Visual rEasoning using top-k generations) that treats base model’s top-k generations, obtained via $K$ rollouts, as training primitive that enables differential visual reasoning. For each training image, we treat the top-k outputs of the base model as an explicit hypotheses set and train the model using RL to resolve this set by selecting the correct element.
We formulate this as a Multiple-Choice-Question (MCQ) interface, using top-k as options leveraging the model’s own distribution. This yields two advantages (i) differential reasoning: presenting the model with its own top predictions as options compels it to move beyond simple pattern recognition. Thus, the model learns to engage in fine-grained reasoning, identifying the specific attributes that differentiate the correct answer from other plausible alternatives, as illustrated in Figure 1. (ii) easily verifiable reward signal: the reward for a correct prediction becomes trivially verifiable as model simply has to select the correct index from the given options. This contrasts with methods such as ViRFT, which rely on an exact string match. Our approach further mitigates the category name memorization issue and leads to better generalization on unseen categories.
Figure 1: For fine-grained image recognition task, most salient visual attributes are often insufficient to identify the correct category as its common among similar categories. (a) This leads to a significant performance gap in model’s Pass@1 and Pass@20 accuracy (b) A differential reasoning can help indicate out the key visual attributes that can help distinguish among similar categories. Base model fails to use such discriminative features relying only on prominent visual features. We solve this by using top-k as options (the most likely categories base model confuses it for) and utilizes model’s text knowledge to resolve this confusion using differential reasoning (highlighted in green).
Experiments on five standard fine-grained image classification datasets show that DiVE-k outperforms existing methods by a significant margin in two distinct zero-shot settings. For standard base-to-novel generalization, our method surpasses pre-trained QWEN2.5-VL-7B and ViRFT by $10.04\%$ and $6.16\%$ on the Harmonic Mean (HM), respectively. This performance gain also extends to mixed-domain zero-shot base-to-novel generalization setting, where we achieve improvements of $9.03\%$ against QWEN2.5-VL-7B and $4.02\%$ against ViRFT. Further, we observe an average improvement of $7.73\%$ compared to ViRFT on 4-shot image classification.
In summary, our main contributions are: (i) we propose DiVE-k framework which uses the top-k generations of base model as a training signal for fine-grained image classification, (ii) we demonstrate the benefits of using MCQ to distinguish among semantically similar categories, and (iii) we show improved performance on multiple fine-grained image classification datasets with detailed ablation studies.
2 Prior Work
Zero-shot fine-grained image classification Vision Language Models (VLMs) (radford2021learning; li2022blip; tschannen2025siglip; yuan2021florence) use the idea of aligning image with text to achieve zero-shot learning (lampert2013attribute; socher2013zero; wang2018zero; zhang2017learning). Although very competent for image-text alignment, these models have limited world knowledge unlike LLMs (radford2019language; brown2020language; touvron2023llama; chowdhery2023palm). An early line of research proposes the idea of prompt learning (zhou2022conditional; khattak2023self) where a prompt vector is learned for text prompt’s context words. zheng2024large uses LLM’s knowledge to learn the prompt vector. Another approach to bridge the knowledge gap is by combining the perceptual strengths of VLMs with the linguistic abilities of LLM (esfandiarpoor2023follow; menon2022visual; pratt2023does; zeng2022socratic; novack2023chils; roth2023waffling). Our work in part is inspired by FuDD (esfandiarpoor2023follow) that uses a multi-stage reasoning by using LLM knowledge to find pairwise discriminative features to compliment VLM. However, FuDD generates a fixed set of text prompts offline, limiting its ability to adapt its reasoning strategy to the specific difficulty of each input. Other related work tries to use LLM as a tool for reasoning through programming and language reasoner (chen2023see; gupta2023visual; zhang2023multimodal; suris2023vipergpt). While these methods improved model performance, the separation between the vision and language modules creates a bottleneck, inhibiting seamless and integrated reasoning across modalities (liu2023visual; liu2024improved).
Recent LVLMs (bai2025qwen2; hurst2024gpt; team2023gemini; team2025gemma) have excellent visual understanding, such as VQA (shao2024visual), combining visual encoding directly into a LLM architecture. This opens up the possibility of joint vision and text reasoning capability (team2025kimi_vl; bai2025qwen2; team2025gemma). Additionally, recent success of RL for reasoning on maths and coding tasks (jaech2024openai; guo2025deepseek; team2025kimi) have transformed the post-training reasoning research and there is a growing interest to extend these ideas to LVLMs.
Enhancing Reasoning with Reinforcement Learning . Building upon the seminal work of In-context learning (brown2020language) and CoT (wei2022chain), the field has moved beyond static prompting by applying RL to fine-tune these reasoning processes (jaech2024openai; guo2025deepseek; team2025kimi). By treating the generation of a reasoning chain as a sequential decision-making problem, RL-based methods can train models to produce more accurate and explainable solutions. Recent breakthrough in DeepSeek-R1 (shao2024deepseekmath) showed the effectiveness of CoT based training further making it more efficient using their GRPO algorithm. Inspired by these success in text-based reasoning, a growing body of work apply RL to enhance the reasoning capabilities of LVLMs for vision-centric tasks, such as image classification (liu2025visual; li2025think), Object detection, Grounding (liu2025visual; shen2025vlm), and Visual Question Answering (cao2025ground; sarch2025grounded; fan2025grit). Within our target domain of fine-grained image recognition, the most pertinent work is ViRFT (liu2025visual), which trains an LVLM using exact string matching reward to foster visual reasoning. However, this reward mechanism proves brittle, failing to generalize in base-to-novel settings. Furthermore, it does not fully leverage the LLM’s inherent knowledge about fine-grained categories to incentivize the differential reasoning necessary for distinguishing among similar options. zhu2025surprising; chen2025pass proposes to enhance LLM performance by improving Pass@k accuracy. In contrast, we use model’s own knowledge of Pass@k to improve their reasoning, specifically for vision-language task.
Figure 2: An overview of DiVE-k framework. First we do an offline option mining (red box) where for each training image, we sample $K$ rollouts from a pretrained LVLM and select top-k options by frequency, ensuring the ground-truth appears. Next we perform RL training using GRPO on MCQ prompts (green box): the model receives an image, a natural language prompt, and k options as input and produces a reasoning chain and a final choice and is optimized with a simple, verifiable reward that combines MCQ correctness and format compliance.
3 DiVE-k Framework
DiVE-k framework employs a simple two step strategy which elicits a differential reasoning in LVLM using its top-k generations as training signal. An overview of our proposed method is shown in Figure 2. In the first step (red box in 2), we perform an offline top-k generation using the base model to construct a potential hypotheses set to be used for constructing Multiple Choice Questions (MCQs). In the second step (green box in 2), we use the MCQ dataset for RL training using GRPO.
top-k as hypotheses set. In the first step of the DiVE-k framework, Given an image $I$ and a text query $q$, we use the base policy model $\pi_{\theta}$ to rollout $K$ responses $\mathcal{{Y}}=({y}_{1},{y}_{2},...,{y}_{K})$ sampled through a specific decoding strategy (e.g. topptop{-}p nucleus sampling).
$\mathcal{{Y}}\sim\pi_{\theta}(I,q,K)$ (1)
Each of these responses ($y_{i}\in\mathcal{{Y}}$) can be represented as $y_{i}=(r_{i},c_{i})$, where $r_{i}$ is the reasoning trace, and $c_{i}$ is the final predicted category name. Let $\mathcal{{C}}$ be the unique category names set within the $K$ generated responses. Next, we count the frequencies of each category in $\mathcal{{C}}$ and using this frequency count, we construct the option set $\mathcal{O}_{top-k}$ by selecting the $k$ most frequent categories. The value of $k$ is set to $k=\min(m,|\mathcal{C}|)$, where $|\mathcal{C}|$ is the number of unique categories, and we use $m=5$. Let $\hat{c}$ be the ground-truth category name. To ensure that the correct answer is always an option during training, we adjust the set if necessary. If $\hat{c}\notin\mathcal{O}_{top-k}$, we modify $\mathcal{O}_{top-k}$ by replacing the least frequent candidate with the ground-truth $\hat{c}$.
[…]
To focus on challenging examples through hard-negative mining, we filter out trivial cases from the training set. Specifically, we exclude any sample for which the model generates only a single, correct category prediction (i.e., $|\mathcal{C}|=1$ and $\mathcal{C}=\{\hat{c}\}$). Note that our first step to generate options with the same base model used as the policy model during training is crucial, as it ensures the categories are drawn from the model’s own distribution and yields optimal learning as shown in section 4.3.1.
Figure 3: An example to illustrate our inference pipeline (red arrows) and its comparison to existing method (blue arrows). Similar to training phase, we perform inference in two steps (right of dotted line), where we first generate option by choosing top-k responses from $K$ rollouts and then model picks the correct answer among the options unlike open-ended one step inference of existing methods (left of dotted line)
[…]
$A_{i}=\frac{r_{i}-\text{mean}\{r_{1},\dots,r_{N}\}}{\text{std}\{r_{1},\dots,r_{N}\}+\delta},$ (2)
where $\delta$ is a small valued constant.
Our reward consists of two parts: first is MCQ reward ($r_{mcq}$) and second is format reward ($r_{format}$). $r_{mcq}$ checks for correctness in the response and rewards a value of 1.0 if model predicts the correct option as answer and 0.0 otherwise as in equation 3
$r_{\mathrm{mcq}}(\hat{a},a)=\begin{cases}1.0,&\text{if }\hat{a}=a\\ 0.0&\text{otherwise.}\end{cases}$ (3)
$r_{format}$ encourages the correct formatting for the output response to make it easier to extract <<think>> and <<answer>> tags. Our final reward ($r$) is defined as the weighted sum of these two: $r=\lambda_{f}r_{format}+\lambda_{m}r_{mcq}$. Our training objective function $\mathcal{J}_{\text{GRPO}}(\theta)$ is defined as:
$\mathcal{J}_{\text{GRPO}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[\min(s_{i}A_{i},\text{clip}(s_{i},1-\epsilon,1+\epsilon)A_{i})-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\right]$ (4)
where $\epsilon$ and $\beta$ are the hyperparameters. $\pi_{\text{ref}}$ is the reference policy model (usually the pre-trained model) used to control the divergence of trained policy. $s_{i}=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$, where $\pi_{\theta_{\text{old}}}$ is the old policy model before the update. Overall, this objective function aims to maximize the expected reward while keeping the policy close to the reference policy for stable learning.
Inference. Figure 3 shows our inference pipeline and compares it to ViRFT which performs a single-pass inference (left of the dotted line) to directly predict the category names, while we use a two-step pipeline (right of the dotted line) similar to training phase. We use the trained policy model to first generate potential options using top-k generations and then re-prompt the same policy to select the correct option among the provided options. It should be noted that we do not add ground-truth to the options if the first step fails to generate the ground-truth category as an option.
4 Experiments
4.1 Experimental Setup
Baselines and Datasets. We compare our results against the various baselines: 1) pre-trained QWEN2.5-VL-7B (bai2025qwen2) 2) Supervised Fine Tuning (SFT) (zheng2024llamafactory) using full model and LoRA (hu2022lora) training. 3) ViRFT (liu2025visual) trained under the same base-novel setting. 4) QWEN2.5-VL-7B and ViRFT using our inference pipeline 5) Consistency as accuracy: During the two-step inference process, use the most consistent prediction from $K$ generation as final prediction. We also report performance of the proprietary models, Gemini2.5-flash-light(comanici2025gemini), GPT-5-mini (gpt5systemcard) and Grok4-fast (grok4fast), with closed weights accessible through APIs (openrouter).
We evaluate our method on five standard fine-grained image classification dataset across various domains, including OxfordFlowers-102 (Nilsback08), CUB-200 (wah2011caltech), OxfordPets-37 (parkhi12a), StanfordCars-196 (krause20133d) and FGVC Aircraft-100 (maji2013fine). For the base-novel split, we follow previous work (Khattak_2023_ICCV) to divide the categories into equal halves for base and novel. For instance in CUB-200, the first 100 categories are considered base and the other 100 are considered novel and similarly for other datasets. Note that neither images nor the category names of the novel classes are seen during training.
Table 1: Quantitative comparison with existing methods on zero-shot base-to-novel generalization. Our proposed methods shows strong generalization outperforming existing methods. B $\rightarrow$ Base, N $\rightarrow$ Novel , H $\rightarrow$ Harmonic Mean, Gemini2.5-f-l $\rightarrow$ Gemini2.5-flash-lite, QWEN2.5 $\rightarrow$ QWEN2.5-VL-7B. Rows with * are results of existing method using our inference pipeline. Consistency $\rightarrow$ Most consistent prediction in top-k generation as final answer.
Flowers CUB Pets Cars Aircraft Avg
Method B N H B N H B N H B N H B N H B N H
Proprietary Models
Gemini2.5-f-l 94.1 91.6 92.8 75.3 60.3 67.0 91.5 96.4 93.9 73.9 92.3 82.1 59.7 65.4 62.4 78.9 81.2 80.0
GPT-5-mini 97.4 94.4 95.9 76.3 64.7 70.0 93.1 97.5 95.2 82.0 94.7 87.9 60.2 74.5 66.6 81.8 85.1 83.4
Grok-4-fast 81.3 87.4 84.2 63.3 50.7 56.3 84.1 94.2 88.8 72.2 86.6 78.7 44.0 55.4 49.1 69.0 74.8 71.8
Method Comparison
CLIP 72.1 77.8 74.8 65.5 48.7 55.8 91.2 97.3 94.1 63.4 74.9 68.6 27.2 36.3 31.1 63.8 67.0 64.9
QWEN2.5 84.2 83.8 84.0 63.3 48.2 54.7 87.5 93.3 90.3 58.9 72.9 65.1 50.3 54.3 52.3 68.9 70.5 69.7
SFT (Full) 93.7 62.6 75.1 84.8 29.2 43.4 95.8 42.6 59.0 81.2 40.6 54.1 64.7 10.1 17.5 84.0 37.0 49.8
SFT (LoRA) 84.5 83.9 84.2 62.2 49.2 54.9 85.7 93.3 89.3 59.9 71.2 65.0 49.5 54.3 51.8 68.3 70.4 69.3
ViRFT 84.3 84.6 84.5 65.4 51.0 57.3 90.5 95.5 92.9 60.3 73.6 66.3 64.6 66.3 65.4 73.0 74.2 73.6
QWEN2.5* 90.7 87.9 89.3 68.5 58.7 63.2 85.6 93.8 89.5 63.3 76.6 69.4 63.3 68.0 65.6 74.3 77.0 75.6
Consistency 85.9 85.5 85.7 64.2 54.0 58.6 90.2 94.7 92.4 62.1 72.0 66.7 61.8 63.5 62.6 72.8 74.0 73.2
ViRFT* 89.5 87.9 88.7 70.0 60.0 64.8 92.6 92.8 92.7 64.8 76.9 70.4 66.4 67.8 67.1 76.8 77.1 76.9
Consistency 85.3 85.2 85.3 67.5 57.2 61.9 92.0 96.1 94.0 64.8 73.2 68.8 67.4 64.2 65.7 75.4 75.2 75.1
DiVE-k (ours) 97.4 88.9 92.9 80.5 65.5 72.2 89.1 94.2 91.6 69.0 76.2 72.4 68.1 69.1 68.6 80.8 78.8 79.8
Consistency 95.3 85.6 90.2 75.2 62.3 68.2 93.1 95.5 94.3 68.7 74.4 71.4 68.0 65.6 66.8 80.0 76.7 78.2
$\Delta$ vs ViRFT +13.1 +4.3 +8.4 +15.1 +14.5 +14.9 -1.4 -1.3 -1.3 +8.7 +2.6 +6.1 +3.5 +2.8 +3.2 +7.8 +4.6 +6.2
$\Delta$ vs QWEN2.5 +13.2 +5.1 +8.9 +17.2 +17.3 +17.5 +1.6 +0.9 +1.3 +10.1 +3.3 +7.3 +17.8 +14.8 +16.3 +11.9 +8.3 +10.1
Evaluation setting and Metric. We evaluate our method under two distinct zero-shot settings. The first is the standard zero-shot base-to-novel generalization, following (Zheng_2024_CVPR), where a separate model is trained on the base classes of each dataset. The second is our proposed mixed-dataset setting, designed to assess cross-domain generalization capabilities. Here, a single model is trained on a unified dataset constructed by combining the base classes from all datasets. Additionally, we also evaluate our method under few-shot classification setting with 4 shots per class. For performance measurement, we report the classification accuracy on base and novel classes, along with their Harmonic Mean (HM).
For evaluation, we use the LLM gemini-2.5-flash-lite (comanici2025gemini), where we provide the ground-truth category name and model predicted category name and ask the LLM if they belong to the same fine-grained category or not. This specifically helps us evaluate better for the answers where model responds a scientific name and the provided ground-truth is common name and vice-versa. We provide more details in Appendix A.1.2
Implementation Details. We use the pre-trained Qwen2.5-VL-7B-Instruct (bai2025qwen2) as our base model and perform the RL training using the proposed method. We use K=20 during offline option generation using $K$ rollouts and other details are provided in Appendix A.1.1. For Zero Shot Base-to-Novel training, we train the model for 400 steps, whereas for mixed data training, we train it for 1 epoch. For few-shot training, we train each model for 200 steps following ViRFT. All models are trained on three A6000 GPUs with 48GB memory with an overall batch size of 6. For GRPO, a total of 4 responses are generated for each input sample. Following ViRFT, we use learning rate of $10^{-6}$, AdamW optimizer, linear scheduler and set $\lambda_{m}=\lambda_{f}=1$. For ViRFT training, we use their official code from github with some modification discussed in Appendix A.1.1.
4.2 Results
4.2.1 Zero-Shot Base-to-Novel Generalization
We present the results for zero-shot base-to-novel generalization in Table 1. Across five fine-grained benchmarks, DiVE-k shows the best generalization, yielding the highest average harmonic mean (HM) of 79.8 with average base/novel accuracies of 80.8/78.8. Relative to the strongest baseline (ViRFT), this corresponds to gains of +7.8 base, +4.6 novel, and +6.2 HM. Even when baselines are run with our inference pipeline (ViRFT*), we retain a +4.0 base, +1.8 novel improvements. Notably, under our inference pipeline, ViRFT gain over QWEN2.5-VL-7B model on novel categories remains marginal by +0.1 (77.0→77.1), suggesting their reward primarily reinforces base categories. In contrast, our approach delivers a substantive +1.8 boost on novel classes under the same inference setting, indicating better generalization beyond the training categories.
The improvements are especially pronounced on CUB (+14.9 HM) and Oxford Flowers (+8.5 HM), and remain consistent on Stanford Cars (+6.1 HM) and FGVC Aircraft (+3.2 HM), indicating robust zero-shot transfer. We observe a small regression on Pet data compared to ViRFT, possibly due to more options leading the model to make more mistake during the second step. We show that our method outperforms ViRFT for smaller $K$ and provide more analysis for this in Ablation 4.3.2.
We also evaluate supervised fine-tuning (SFT) as a baseline. Although SFT yields strong accuracy gains on base categories, its performance deteriorates sharply on novel categories, with an average accuracy drop of 33.5% relative to the base model and a 19.9% reduction in HM. This sharp degradation highlights SFT’s inability to generalize under the base-to-novel transfer setting. To provide a broader perspective, we also include results from proprietary models which are likely much larger. In this context, our method surpasses Grok4-fast and is on par with Gemini2.5-flash-light, though GPT-5-mini leads. We provide more results using Gemma-3 (team2025gemma) base model in Appendix A.3.
Table 2: Quantitative comparison of our method with baselines under mixed-dataset base-to-novel generalization. QWEN2.5 $\rightarrow$ QWEN2.5-VL-7B. Rows with * are results of existing method using our inference pipeline. Consistency $\rightarrow$ Most consistent prediction in top-k generations as output.
Flowers CUB Pets Cars Aircraft Avg
Method B N H B N H B N H B N H B N H B N H
QWEN2.5 84.2 83.8 84.0 63.3 48.2 54.7 87.5 93.3 90.3 58.9 72.9 65.1 50.3 54.3 52.3 68.9 70.5 69.7
ViRFT 87.2 85.9 86.5 65.7 54.0 59.3 90.5 95.0 92.7 59.7 73.6 66.0 67.2 68.0 67.5 74.1 75.3 74.7
QWEN2.5* 90.7 87.9 89.3 68.5 58.7 63.2 85.6 93.8 89.5 63.3 76.6 69.4 63.3 68.0 65.6 74.3 77.0 75.6
Consistency 85.9 85.5 85.7 64.2 54.0 58.6 90.2 94.7 92.4 62.1 72.0 66.7 61.8 63.5 62.6 61.9 67.8 64.6
ViRFT* 90.2 87.9 89.0 71.3 58.3 64.2 90.2 91.6 90.9 63.7 76.6 69.6 66.9 66.4 66.7 76.5 76.2 76.1
Consistency 85.2 85.1 85.2 68.0 57.5 62.3 91.5 96.1 93.7 65.6 74.2 69.7 63.2 65.5 64.3 64.4 69.9 67.0
DiVE-k (ours) 97.4 89.9 93.5 76.8 61.3 68.2 87.8 94.7 91.1 68.5 78.5 73.1 65.5 69.7 67.5 79.2 78.8 78.7
Consistency 92.8 86.0 89.3 67.7 56.7 61.7 91.8 96.1 93.9 59.6 72.5 65.4 61.1 63.5 62.3 60.4 68.0 63.9
$\Delta$ vs ViRFT +10.2 +4.0 +7.0 +11.1 +7.3 +8.9 -2.7 -0.3 -1.6 +8.8 +4.9 +7.1 -1.7 +1.7 0.0 +5.1 +3.5 +4.0
$\Delta$ vs QWEN +13.2 +6.1 +9.5 +13.5 +13.1 +13.5 +0.3 +1.4 +0.8 +9.6 +5.6 +8.0 +15.2 +15.4 +15.2 +10.3 +8.3 +9.0
4.2.2 Mixed-Dataset Base-to-Novel Generalization
Training a single model on the union of base categories from all five datasets provide a strong evaluation of mixed-dataset generalization. Table 2 shows the quantitative results for this setting. Our method attains the highest average harmonic mean (HM) of 78.7, improving over the pretrained QWEN2.5‑VL‑7B by +9.0 HM and over ViRFT by +4.0 HM, with average base/novel gains of +5.1/+3.5. Under our two-step inference on novel classes, ViRFT underperforms the pretrained QWEN2.5‑VL‑7B (76.2 vs. 77.0), while our approach reaches 78.8, a +1.8 improvement. This contrast suggests that ViRFT struggles to transfer when trained on the mixed base corpus, likely reinforcing base-only cues, whereas our method maintains robust generalization to unseen categories.
Table 3: Quantitative comparison of our proposed method under 4-shot setting.
Model Oxford Flowers CUB Oxford Pets Stanford Cars FGVC Aircraft Average
QWEN2.5-VL-7B 78.43 51.62 79.05 57.91 52.50 63.90
ViRFT 81.12 51.75 85.81 57.65 58.75 67.02
QWEN2.5-VL-7B* 85.04 56.87 80.40 63.46 64.67 70.09
   consistency 81.86 53.00 85.81 58.30 56.36 67.07
ViRFT* 84.80 60.00 83.10 64.75 65.44 71.62
   consistency 83.33 54.12 87.83 59.02 54.71 67.80
DiVE-k (Ours) 88.72 63.87 85.14 66.90 69.10 74.75
   consistency 85.78 58.00 86.48 60.17 64.92 71.07
$\Delta$ w.r.t ViRFT 7.60 12.12 -0.67 9.25 10.35 7.73
$\Delta$ w.r.t QWEN2.5-VL-7B 10.29 12.25 6.09 8.99 16.60 10.85
4.2.3 Few-Shot Classification
DiVE-k also shows improvement under few-shot classification setting across datasets. As shown in Table 3, our method achieves an average HM classification accuracy of 74.75%, an improvement of 7.73% compared to ViRFT and 10.85% compared to QWEN2.5-VL-7B model, demonstrating the effectiveness of our method even under data efficient training.
Figure 4: Qualitative comparison on fine-grained flower recognition (top: ViRFT; bottom: Ours). Top: ViRFT predicts “global thistle,” which is incorrect and reflects a coarse judgment. Bottom: Our method enumerates close candidates and uses attribute-grounded, differential reasoning such as capitulum/head shape, floret density and arrangement, bract patterning to select the correct fine-grained label with a justification aligned to the final choice.
4.2.4 Visualization
In Figure 4, we visualize and compare DiVE-k to ViRFT and defer additional visualizations to Appendix A.2. We find that ViRFT directly latches onto high-level “thistle-like” cues and commits to an incorrect category (“global thistle”), without checking the discriminative attributes that separate near-neighbors. In contrast, DiVE-k first proposes a small top-k shortlist and then rules candidates in/out through explicit, attribute-level comparisons, such as head geometry, the density/arrangement of florets, and the presence/shape of bracts, before committing to a final option. This process yields the correct species-level label together with a rationale that stays consistent with the answer reducing overgeneralization and improving interpretability.
4.3 Ablation Studies
4.3.1 The Efficacy of top-k for MCQ Option Generation.
Table 4: Quantitative comparison of classification accuracy on CUB dataset when options are sampled in different ways.
Base Novel HM
QWEN2.5-VL-7B 68.5 58.7 63.2
Random MCQ 72.5 57.5 64.1
Text Emb MCQ 76.0 59.3 66.8
top-k MCQ 80.5 65.5 72.2
Our ablation into the MCQ option generation reveals that the option construction strategy is critical for model’s performance. As detailed in Table 3, randomly selecting categories proved suboptimal, yielding only marginal gains and failing to instill robust reasoning capabilities. While employing a text embedding model from gemini embeddings (lee2025gemini) to generate semantically similar options offered some improvement, our proposed top-k as options proves significantly more effective. By sampling options directly from the same base model’s top-k generations achieved a substantial classification accuracy gain of 12% on base classes and 6.9% on novel classes compared to base model. This demonstrates that leveraging the base model’s own knowledge distribution to generate options is the optimal strategy for training, leading to superior generalization on both base and novel sets.
4.3.2 Role of top-k generations During Inference
Figure 5: Change in classification accuracy for different values of $K$ on different dataset. Our approach consistently outperforms the baselines for nearly all $K$.
We analyze the impact of the hyperparameter $K$, during $K$ rollout for top-k generation, on classification accuracy, with results presented in Figure 5. The plots show the HM of accuracy across base and novel split for different values of $K$ during inference. Our method consistently surpasses both the baselines at nearly all $K$, often by a large margin. As $K$ increases, accuracy generally rises and then saturates around $K\sim 10-15$, yielding near-maximal performance at K=1520K=15-20 for Flowers ($\sim$93%), CUB ($\sim$72.5%), Car ($\sim$73%), and Aircraft ($\sim$70%), indicating diminishing returns beyond K=1015K=10–15. However, for Pet dataset it peaks at small $K$ (93.8% at $K=2$) and gradually declines for larger $K$, suggesting more options during MCQ leads to more mistakes for this specific dataset. Overall, these results show that our approach achieves improved performance even for small $K$ and thus can get benefit while avoiding extra computation.
4.3.3 Roles of Vision and Text Components
Table 5: Ablation on effect of training different part of the model on classification accuracy for CUB dataset; Full model training yields the best performance.
Base Novel HM
QWEN2.5-VL-7B 68.50 58.67 63.21
Vision only training 74.00 57.83 64.92
Text only training 74.33 60.33 66.60
Full model training 80.50 65.50 72.23
To investigate the individual contributions of the model’s vision and text components to its reasoning capabilities, we conduct an ablation to determine whether performance gains stem primarily from updating the vision features, refining the language model’s ability to generate a correct chain of reasoning tokens, or a combination of both. the results of which are presented in Table 5.
Our findings reveal distinct roles for each modality. When we fine-tuned only the vision components (freeze text decoder), the model’s performance improved significantly on the base dataset (68.5→74) but failed to generalize to the novel dataset, where performance slightly degraded (58.67→57.83). This suggests that adapting visual features alone is insufficient for robust reasoning on new, unseen data. Conversely, training only the text components (freeze vision tower) improved performance on both the base (68.50→74.33) and novel (58.67→60.33) sets. This indicates that enhancing the language model’s ability to generate logical reasoning is critical for generalization. However, the best performance was achieved through full model training, which yielded substantial gains on both base (68.50→80.50) and novel (58.67→65.50) sets. This demonstrates that while language model adaptation is key, a combination of both vision and text modules is necessary to unlock the model’s full reasoning capability.
4.3.4 Accuracy Analysis of the Two Steps in DiVE-k
As DiVE-k framework operates in two stages, the second stage can only succeed if the correct answer appears in the top-k rollouts from the first stage. Therefore, understanding the bottleneck between these two steps is crucial for diagnosing model performance.
Table 6 reports the top-k accuracy (Step 1) of both the base QWEN2.5-VL model and the DiVE-k–trained model. We find that the base model already achieves strong top-k recall across most datasets, indicating that the correct answer is typically recoverable via sampling. DiVE-k training further improves this recall, yielding more consistent retrieval of the correct candidates. This improvement is particularly beneficial because Step 2 can only operate correctly when Step 1 supplies the correct option.
Table 6: Top-k accuracy (step 1) across five datasets
Flowers CUB Pets Cars Aircraft Avg
Method B N B N B N B N B N B N
QWEN2.5-VL 96.84 92.62 83.33 74.66 98.40 98.89 86.67 90.84 89.02 90.15 90.9 89.4
DiVE-k (ours) 98.99 92.62 89.66 77.50 98.94 99.16 89.74 90.57 92.96 90.21 94.1 90.0
$\Delta$ 2.15 0.00 6.33 2.84 0.54 0.27 3.07 -0.27 3.94 0.06 3.2 0.6
Table 7 presents the Step 2 MCQ accuracy (differential reasoning). We observe an average improvement of 4.3% on base and 1.6% on novel categories. Since Step 2 operates over the candidates generated in Step 1, high top-k recall directly strengthens the effectiveness of differential reasoning. Improvements in candidate quality and reasoning accuracy reinforce each other, resulting in a compounding effect on the final classification performance.
Across most datasets, Step 1 accuracy is already high (often above 90%), which shifts the primary performance bottleneck to the differential reasoning stage. DiVE-k explicitly targets this challenge and achieves clear gains in MCQ accuracy. For the CUB dataset, however, top-k accuracy remains comparatively lower, leaving additional room for improvement in Step 1. This highlights that the two stages contribute differently depending on dataset difficulty. Some benchmarks are limited by candidate generation, while others are limited by reasoning over those candidates.
Table 7: MCQ accuracy (Step 2) across five datasets
Flowers CUB Pets Cars Aircraft Avg
Method B N B N B N B N B N B N
QWEN2.5-VL 93.65 94.90 82.20 78.62 86.99 94.85 73.03 84.32 71.10 75.42 81.4 85.6
DiVE-k (ours) 98.40 95.98 89.78 84.52 90.05 95.00 76.90 84.10 73.25 76.59 85.7 87.2
$\Delta$ 4.75 1.08 7.58 5.90 3.06 0.15 3.87 -0.22 2.15 1.17 4.3 1.6
5 Limitations and future work
While DiVE-k framework effectively leverages the model’s intrinsic knowledge acquired during pre-training to improve its downstream accuracy, our two-step inference process incurs additional computational cost due to the requirement of two forward passes. The success of our approach is demonstrated on QWEN2.5-VL, which exhibits high initial Pass@k accuracy. However, the method’s efficacy is contingent on this baseline performance; base LVLMs with lower intrinsic accuracy may not realize comparable gains. A potential direction to mitigate this dependency, which we leave for future work, is to incorporate a Pass@k accuracy as reward signal directly into the training objective. Another promising avenue for future research involves the verification of generated reasoning traces for factual correctness and their grounding in the input image.
6 conclusion
In conclusion, we introduced DiVE-k, a novel framework that addresses the limitations of Large Vision Language Models in fine-grained image recognition. By utilizing top-k generations as training primitive, our method requires the model to perform differential reasoning among visually similar categories using a multiple-choice question format. Extensive experiments across five standard datasets demonstrate that DiVE-k significantly outperforms existing approaches in base-to-novel generalization, mixed domain, and few-shot settings. Our ablation studies further reveal that the efficacy of this approach hinges on mining options from the base model’s own distribution, which is critical for effective RL training. Moreover, we show that the joint fine-tuning of both vision and text components is essential for unlocking the model’s full reasoning potential and that increasing the value of k offers diminishing return during inference. Overall, our work highlights the effectiveness of leveraging a model’s inherent knowledge distribution to refine its reasoning capabilities, establishing a new direction for improving visual discrimination in LVLMs.
7 Reproducibility statement
We aim to ensure full reproducibility of our work by providing detailed descriptions of our methodology, training and inference pipeline, and evaluation protocol in the method, results and appendix section. All hyperparameters, implementation details, and generation settings (e.g., temperature, sampling strategies, and reward design) are listed in the Appendix, along with the exact prompt templates used for all experiments. Experiments are conducted on both open-source and closed-source models; for open-source models, we provide precise checkpoint versions. We plan to publicly release our codebase, experimental configurations, and trained model checkpoints upon acceptance to support reproducibility.‘
Appendix A appendix
A.1 Implementation Details
A.1.1 Training details
Sampling parameters. In the first step of our pipeline, we generate $K$ responses using top-p nucleus sampling. We provide the details about the parameters using during this sampling in Table 8.
Table 8: Generation Arguments
Parameter Value
max_new_tokens 1024
temperature 1.0
top_p 0.95
do_sample True
num_return_sequences 20
repetition_penalty 1.1
ViRFT model training. During our evaluation of ViRFT method, we find a crucial issue in their implementation. The string match logic used during both training (for reward computation) and evaluation (for answer correctness) is shown in Listing A.1.1. We find that the second part of the or statement (student_answer in ground_truth) leads to a shortcut specifically for fine-grained image classification. For example, even if the model responds “gull” this reward function (and evaluation) will consider it as correct even though it doesn’t give a correct answer. (it could be any of “california gull”, “Heermann Gull”, “ivory gull” etc.). This leads to a huge drop in accuracy when evaluated using our LLM evaluation as demonstrated in Table 9 for CUB and Stanford Cars dataset. We fix this issue by remove this shortcut and keeping only the “ground_truth in student_answer” during training for reward computation. After fixing this error we observe expected results shown in 9.
Figure 6: Qualitative comparison across three domains—aircraft, birds, and cars. In each pair, ViRFT commits to a plausible but incorrect class with generic rationale such as “BAe 146-300”, “Great Crested Flycatcher”, “2012 Ford F-150”. Our method first enumerates top-k candidates and then applies attribute-grounded differential reasoning such as T-tail/registration cues for BAe 146-200; Empidonax traits for Least Flycatcher; grille/headlight era cues for a 2007 F-150, yielding the correct fine-grained label and a justification aligned with the final choice.
Table 9: Table to demonstrate the shortcut issue in original ViRFT code. ViRFT refers to the accuracy using the original code, ViRFT! refers to our modified code used for training
Method CUB Stanford cars
Base Novel Base Novel
QWEN2.5-VL-7B 63.33 48.17 58.87 72.9
ViRFT 39.33 27.83 13.79 23.09
ViRFT! 65.44 51.00 60.34 73.63
String matching code in ViRFT 1# reward computation code 2if ground_truth in student_answer or student_answer in ground_truth: 3 reward = 1.0 4 5# evaluation code 6if image_cate in answer_content or answer_content in image_cate: 7 right_count += 1 8else: 9 print(’no’)
<image> This is an image containing a bird. Output the most likely species name in the image. The species name of the bird strictly belongs to below category list {category_list}. Output the thinking process in <think> </think> and final answer in <answer> </answer> tags. The output answer format should be as follows: <think></think> <answer> species name </answer>. Please strictly follow the format.
Figure 7: Prompt for first step of our pipeline to generate $K$ rollouts
This is an image containing a bird. Please find the most likely bird in the image from the below options. {options}. Please output the letter corresponding to the correct category name. Output the thinking process in <think> </think> and final answer in <answer> </answer> tags. The output answer format should be as follows: <think></think> <answer>option letter</answer>
Please strictly follow the format.
Figure 8: Prompt for Multiple Choice Question (MCQ) answering
You are evaluating fine-grained image classification results. Given: - Groundtruth category: {groundtruth} - LLM prediction: {prediction} Check if the groundtruth matches the prediction. The strings need not match exactly but they must refer to the same specific fine-grained category, not just broad class. Respond with: 1. ”True” or ”False” if groundtruth matches the prediction in <answer></answer> tag. i.e <answer>answer here (True/False)</answer>
2. Brief explanation in <explanation></explanation> tag. i.e <explanation>Explanation here</explanation>
Figure 9: Prompt for evaluating fine-grained image classification results.
A.1.2 Evaluation details
Here we provide the details of the different prompt used at different stage of our method. In Figure 7 we have provided the complete prompt used in first step of our proposed method. Here, {category_list} refers to the list of all the category list for the given dataset. During training, we only use base categories in the prompt.
In Figure 8, we provide the prompt used for second step of multiple choice question. Here {options} refers to the options obtained from the first step of our pipeline. In Figure 9, we provide the prompt used during the evaluation. {groundtruth} refers to the groundtruth category name and {prediction} refers to the model’s predicted answer. We use "google/gemini-2.5-flash-lite-preview-06-17" as LLM for evaluation from openrouter (openrouter) API.
A.2 Qualitative comparison
We provide additional qualitative comparisons in Figure 6. We note that explicit candidate enumeration followed by differential, attribute-level reasoning improves fine-grained recognition. In each case, the top rows (ViRFT) select look-alike but wrong categories—over-generalizing to BAe 146-300, misidentifying a flycatcher species, and over-estimating the truck’s model year—supported by broad, non-discriminative explanations. The bottom rows (Ours) surface a short top-k list and then contrast salient cues (e.g., tail/engine/registration details for BAe 146-200; size/underparts/Empidonax patterns for Least Flycatcher; grille and headlight silhouette for a 2007 F-150) before committing to a final answer. This two-step structure reduces over-generalization and aligns the selected label with evidence visible in the image.
Table 10: Comparison of classification accuracy using Gemma3-12B base model on CUB and Flowers datasets.
CUB Flowers
Method Base Novel HM Base Novel HM
Gemma3-12B 38.67 32.00 35.03 71.26 86.52 78.16
Gemma3-12B* 46.00 37.17 41.12 69.54 86.67 77.17
DiVE-k (Ours) 58.50 40.83 48.10 84.63 87.38 85.99
$\Delta$ w.r.t Gemma3 20.17 8.83 13.07 13.37 0.86 7.83
A.3 Additional Backbone Results
To assess the model-agnostic nature of our approach, we also report results with replacing the QWEN2.5-VL backbone with Gemma3-12B and repeat the evaluation on CUB and Flowers. As shown in Table 10, DiVE-k consistently yields substantial improvements over the Gemma3-12B baseline across all splits.Overall, DiVE-k consistently improves the harmonic mean (HM) performance on both datasets, achieving gains of +13.07 and +7.83 points over the base Gemma3-12B model. These results demonstrate that our framework generalizes beyond a specific foundation model and delivers consistent performance gains across different architectures and datasets.
A.4 Computation Cost Comparison
Table 11: Per-sample inference time comparison between one-step and two-step pipelines.
One Step Two Steps
Flowers 2.50 s 16.02 s
CUB 2.77 s 14.58 s
Pet 2.23 s 8.25 s
Average 2.50 s 12.95 s
Since DiVE-k uses a two-step inference pipeline, it incurs additional computational overhead due to two forward passes. To quantify this overhead, we measure the per-sample inference time averaged over 500 samples on A6000 GPUs (48GB). As shown in Table 11, the average per-sample inference time increases from 2.50s in the one-step setting to 12.95s in the two-step setting. However, these improvements are not merely a byproduct of increased computation. Even under an identical compute budget, using K=1, which reduces our method to a single forward pass with greedy decoding, directly comparable to prior one-step baselines, DiVE-k still outperforms existing methods on 4 out of 5 datasets (Figure 5). This demonstrates that the gains arise from our formulation itself rather than additional compute alone. When more computation is permitted, performance further scales with K, providing a controllable trade-off between inference cost and accuracy that can be adapted to different application constraints.
A.5 Large Language Models (LLMs) usage detail.
We utilized LLMs as a writing aid. Their application was strictly limited to proofreading for errors and polishing the prose for clarity and style. LLMs were not used for any substantive tasks, including but not limited to research, information retrieval, discovery, or the ideation of concepts and conclusions presented herein. All intellectual content is the original work of the authors.